OpenStudy (anonymous):

A person is watching a boat from the top of a lighthouse. The boat is approaching the lighthouse directly. When first noticed the angle of depression to the boat is 18°33'. When the boat stops, the angle of depression is 51°33'. The lighthouse is 200 feet tall. How far did the boat travel from when it was first noticed until it stopped? Round your answer to the hundredths place.

OpenStudy (anonymous):

@hartnn

OpenStudy (dumbcow):

 we want to find distance "x" diagram creates 2 right triangles , we dont know the hypotenuse so use tangent \[\tan (51'33'') = \frac{200}{y}\] \[\tan (18'33'') = \frac{200}{x+y}\]

OpenStudy (anonymous):

okay..

OpenStudy (dumbcow):

having computer issues :{ ok now using substitution you can solve those 2 equations for "x" \[y = \frac{200}{\tan(51'33'')}\] and \[x = \frac{200}{\tan(18'33'')}-y\] so \[x = \frac{200}{\tan(18'33'')}-\frac{200}{\tan(51'33'')}\]

OpenStudy (anonymous):

okay

OpenStudy (dumbcow):

haha did you follow what i did at all ?

OpenStudy (anonymous):

to be honest , no i did not .